OpenStudy (anonymous):

What is the y–intercept of the graph of 3x + 4y = 12 ?

OpenStudy (anonymous):

y-intercept is when x = 0 3 (0) + 4y = 12 4y = 12 y = 3 y-intercept is (0, 3)

OpenStudy (anonymous):

The equation has the form Ax +By = C In this equation, A=3, B=4 and C=12 We need to put it in the form y+Dx=E. To do this we need to divide through by the coeffecient of y (4 in this case) This would give us y + 0.75x = 3 The constant term is the y intercept So the y intercept of this line is 3 (this occurs when x=0)

OpenStudy (anonymous):

Or you can do it the way moneybird did. It's easier.